A. List the members of the following sets

1. {x| x is real numbers and x2 = 1}

2. {x| x is an integer and -4 < x ≤ 3}

B. Use set builder notation to give description of each of these sets.

1. {a, e,i ,o, u}

2. {=2, -1, 0, 1, 2}

C. Let A= (a, b, c), B = (x, y) and C = (0, 1)

Find:

1. A U C

2. C x B

3. B – A

4. (A ∩ C) U B

D. Find these terms of the sequence (An}, where An = 2(3)n + 5

1. A0

2. A5

3. A3

4. 8th term

5. 2nd term

6. Sum of the sequence 

E. Given the following set:

2. X = {-1, 0, 1, 2, 3, 4, 5} defined by the rule (x, y) ∈R if x ≤ y

F. List the elements of R

G. Find the domain of R

H. Find the range of R

I. Draw the digraph

J. Properties of the Relation 


1
Expert's answer
2021-03-16T11:30:08-0400

Solution.

A.

"\{x| x \text{ is real numbers and } x^2=1\}=\{-1,1\}."

"\{x| x \text{ is an integer and } -4<x\leq 3\}=\{-3,-2,-1,0,1,2,3\}."

B.

"\{a,e,i,o,u\}=\{x| x\text{ is vowels letters, without 'y' }\}."

"\{-2,-1,0,1,2\}=\{x \in Z| |x|=2\}."

C.

"A=(a,b,c), B=(x,y), C=(0,1)."

"A \text{U} C=\{a,b,c,0,1\}. \newline
C \times B=\{(0,x),(0,y),(1,x),(1,y)\}. \newline
B-A=\{x,y\}. \newline
(A \cap C)\text{U} B=\{x,y\}."

D.

"A_n=2\cdot 3^n+5. \newline
A_0=7. \newline
A_5=491. \newline
A_3=59. \newline
A_8=13127.
\newline
A_2=23. \newline
\sum A_n=\infty."

E.

"X=\{-1,0,1,2,3,4,5\}. \newline
R=\{(x,y)|x\leq y\}."

F.

"R=\{(-1,0),(-1,1),(-1,2),(-1,3),(-1,4),(-1,5),\newline
(0,1),(0,2),(0,3),(0,4),(0,5),(1,2),(1,3),(1,4),(1,5),(2,3),(2,4),\newline
(2,5),(3,4),(3,5),(4,5),(5,5),(4,4),(3,3),(2,2),(1,1),(0,0),(-1,-1)\}."

G.

Domain("R" )="\{-1,0,1,2,3,4,5\}."

H.

Range("R" )="\{-1,0,1,2,3,4,5\}."

I.

J.

Properties of the Relation:

Reflexivity

Irreflexivity

Symmetry

Antysymmetry

Asymmetry

Transitivity.